What is the pH of a solution in which 0.0100 grams of HCIO4 is dissolved in 1 liter of solution?
Answer #1

Molar mass of HClO4 = 100.46 g/mol

Moles of HClO4 = 0.01/100.46 = 9.95×10-5 moles

Concentration of H+ = moles/volume (L)

= 9.95×10-5/1 = 9.95×10-5 M

pH = -log[H+] = -log (9.95×10-5) = 4